OpenStudy (anonymous):

I WILL FAN AND MEDAL!!!!!!!!!!!!!!!!!! In the equation y equals one-fourth x plus 8, what is y when x is twenty? Choose one answer. A. twenty-eight B. thirteen C. twelve D. 7

OpenStudy (anonymous):

@ExoticInspirit @elxation

OpenStudy (anonymous):

B :) Do you want me to explain it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@elxation

OpenStudy (anonymous):

Ok so your original equation is \[Y=\frac{ 1 }{ 4 }x +8\]

OpenStudy (anonymous):

First, we plug in 20 for x \[Y=\frac{ 1 }{ 4 }(20)+8\]

OpenStudy (anonymous):

Since all whole numbers have a denominator of 1, we can multiply them together.\[\frac{ 1 }{ 4 } \times \frac{ 20 }{ 1 }\] To do this, you can simply just multiply 1 by 20, and 4 by 1

OpenStudy (anonymous):

Now we get \[Y= \frac{ 20 }{ 4 }+8\] To finish solving, do \[20\div4=5\] and then add 8. 8+5=13